OpenStudy (anonymous):

(help!!) Solve the compound inequality. 36 ≥ 1-5z > -21

OpenStudy (oaktree):

Start off by clearing off the 1...subtract 1 from all sides, leaving 35<-5z<-22

OpenStudy (oaktree):

Now, divide by -5 and get -7>z>22/5

OpenStudy (anonymous):

Split the compound inequality into two inequalities: \[36\ge(1-5z)\] and \[(1-5z)>-21\]

OpenStudy (anonymous):

solve each inequality separately \[36\ge(1-5z)\] subtract one from each side, then divide by (-5) from each side and flip the direction of the inequality sign since we are dividing by a negative. \[35\ge(-5z)\] \[-7\le z\]

OpenStudy (anonymous):

Do the same for the other inequality subtract one and then divide by (-5) and switch the direction of the inequality sign \[(1-5z)>-21\] \[(-5z)>-22\] \[z<\frac{ 22 }{ 5 }\]

OpenStudy (anonymous):

just realized the original process given by OakTree is fine except that it started with the inequality signs pointing the wrong way. \[-7\le z < \frac{ 22 }{ 5 }\]

OpenStudy (anonymous):

@juantweaver Ohhhh thank you so much!!! :) Really appreciate it. I see now :) I wasn't sure to leave it as a fraction or if i needed to subtract the one but now I understand, thanks :)

OpenStudy (anonymous):

& You're the best for showing all the steps so I could understand! Thank you :) I was on the right track!